# Large Language Models Are Cross-Lingual Knowledge-Free Reasoners

### v1

@@ Appendix G Additional Results of Experiment
Figure 12: Left: Accuracy of different models on StrategyQA. Solid and dashed line represent the result of With Facts and No Facts setting, respectively. Middle: Accuracy of different models on StrategyQA before fine-tuning. Right: Accuracy of LLaMA-2-7B-Chat on StrategyQA under various settings. The translucent line represents the accuracy before finetuning on the specific tasks (which are all around 50%).
@@
Figure 13: Accuracy of various models on different parts of KFRD. The translucent line represents the accuracy before finetuning on the specific tasks.
@@
Figure 14: Averaged Accuracy on English and Arabic/Hebrew KFRD for models in different stages trained in Arabic/Hebrew

### v2

@@ Appendix G Reasons for Creating a New Dataset
Figure A2: Left: Accuracy of different models on StrategyQA. Solid and dashed line represent the result of With Facts and No Facts setting, respectively. Middle: Accuracy of different models on StrategyQA before fine-tuning. Right: Accuracy of LLaMA-2-7B-Chat on StrategyQA under various settings. The translucent line represents the accuracy before finetuning on the specific tasks (which are all around 50%).
Figure A3: XLTR of different models on QASC. Solid lines: WF-2 results; Dashed lines: NF results.
Figure A4: XLTR of LLaMA-2-7B-Chat on QASC. Here WF-2 equals to WF-all, as QASC only has two pieces of evidence per sample.
@@
Figure A5: Accuracy of various models on different parts of KFRD. The translucent line represents the accuracy before finetuning on the specific tasks.
Figure A6: Accuracy of LLaMA-2-7B-Chat on three parts of KFRD. The different lines indicate different trained languages.
@@
Figure A7: Averaged Accuracy on English and Arabic/Hebrew KFRD for models in different stages trained in Arabic/Hebrew
MKQA
Query Who sings "I Hear You Knocking But You Can’t Come In"?
Answers Dave Edmunds
BoolQ
Question Do Iran and Afghanistan speak the same language?
Answer True
AmbigQA
Answer 74 days
Table A8: Examples of adapted datasets used in this paper
Appendix H Additional Works on Cross-lingual Transfer
Devlin (2018) introduced mBERT, advancing cross-lingual transfer by capturing shared linguistic patterns in a unified embedding space, enabling zero-shot transfer without parallel corpora. Similarly, Conneau (2019) showed XLM’s effectiveness in optimizing multilingual embeddings, improving performance in translation and classification tasks. Ansell et al. (2021) proposed composable sparse fine-tuning, selectively fine-tuning sparse parameters across languages to reduce interference and boost performance, especially in low-resource settings, outperforming adapter-based methods in tasks like NER and NLI.
Additionally, Ye et al. (2023) assessed the imbalance of knowledge across different languages in LLMs, observing weak cross-lingual transferability of knowledge. Zhu et al. (2024) discovered that training on translated questions can enhance the cross-lingual transferability of reasoning tasks.